Find the amount of an annuity with $2500 deposited quarterly at 8% for four years

A. $46,598.21
B $3,300.00
C. $13,200.00
D $43,200.00

Answer by stanbon(75887)

Find the amount of an annuity with $2500 deposited quarterly at 8% for four years

A. $46,598.21
B $3,300.00
C. $13,200.00
D $43,200.00
---------------
Formula: S = R[(1+i)^n - 1]/i
S is Future Value
R is Periodic deposit
i=0.08/4 = 0.02 is periodic rate.
n= 4*4 is number of periods.
-----------------
S = 2500[1.02^(4*4) - 1]/[0.08/4]
S = 2500[.3727857/0.02]
S = 2500[18.639285]
S = $46,598.21
=============
Cheers,
Stan H.